SPIN-ROTATION INTERACTIONS IN CHLOROMETHYL RADICAL

\underline{A. C. BLAIR}, T. HÄBER, and D. J. NESBITT, JILA, University of Colorado, Boulder, Colorado 80309-0440.

Velocity collimation in the slit supersonic discharge expansion results in a strong reduction in residual Doppler broadening, which makes possible resolution of spin-rotation fine structure in open shell radicals in the near infrared. By way of example, spin-rotation splittings in the vibrational ground and excited state have been observed in the ro-vibrational spectrum of the symmetric C-H-vibration (3055.07723(63) cm^{−1})of chloromethyl radical, for both K = 0 and 1 nuclear spin species. Rovibrational line shapes have been fitted to a spin-rotation Hamiltonian, including rotational-asymmetry splitting as well as hyperfine interactions. Spin-rotation coupling constants have been obtained for the vibrational excited state, with ground state values taken from microwave spectra.^{a}

aY. Endo, S. Saito, E. Hirota, Can. J. Phys. 62, 1347 (1984)